OpenStudy (allieatthedisco):

If y varies directly as x, and y=10 when x=2.4, find x when y=15.

jimthompson5910 (jim_thompson5910):

Since "y varies directly as x", we know that y = k*x where k is some fixed number

jimthompson5910 (jim_thompson5910):

we are told that "y=10 when x=2.4", so plug these values into the equation to get y = k*x 10 = k*2.4 now solve for k and tell me what you get

OpenStudy (allieatthedisco):

would k = 4.16?

jimthompson5910 (jim_thompson5910):

roughly, yes

jimthompson5910 (jim_thompson5910):

so your equation y = k*x updates to y = 4.16*x (approximately)

jimthompson5910 (jim_thompson5910):

once you know k, you can find x when given a certain value of y (or vice versa)

jimthompson5910 (jim_thompson5910):

this time we need to find x when y = 15 y = 4.16*x 15 = 4.16*x x = ??

OpenStudy (allieatthedisco):

x=3.60

jimthompson5910 (jim_thompson5910):

good, that is the correct answer

OpenStudy (allieatthedisco):

ok thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome